Chapter 10 Homework A Saved On January 1, 2018, Eagle borrows $15,000 cash by signing a four-year, 7% installment note. The nTABLE B.1 Present Value of 1 p=1/(1+i) Rate Periods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% 0 0 0.9901 0.9803 0.9706 0.9610TABLE B.2 Future Value of 1 f= (1 + i) Rate Perlods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% 1.0000 1.0100 1.0201 1.0303 1.040TABLE B.3 Present Value of an Annuity of 1 Rate Perlods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% No von WN 0.9901 1.9704 2.9410f=[(1 + i) – 19i TABLE B.45 Future Value of an Annuity of 1 Rate Periods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% O VOWN 1.000
Answer #1
Note Repayment Schedule
Year Opening balance Interest Installment Principal Repayment Closing Balance
a b c d e=d-c f=b+c-d
2018 $                                            15,000 $          1,050 $          4,428 $                             3,378 $        11,622
2019 $                                            11,622 $             813 $          4,428 $                             3,615 $          8,007
2020 $                                              8,007 $             560 $          4,428 $                             3,868 $          4,139
2021 $                                              4,139 $             289 $          4,428 $                             4,139 $                 -  
Journal Entries
Date General Journal Debit Credit
January 1,2018 Cash $       15,000
Note Payable $       15,000
(to record borrow of money)
December31,2018 Interest Expenses 1050
Note Payable 3378
cash 4428
(repayment of interest & principal)
December31,2019 Interest Expenses 813
Note Payable 3615
cash 4428
(repayment of interest & principal)
December31,2020 Interest Expenses 560
Note Payable 3868
cash 4428
(repayment of interest & principal)
December31,2021 Interest Expenses 289
Note Payable 4139
cash 4428
(repayment of interest & principal)